A firm produces candles. The market for candles is highly competitive, with candles currently selling for $10. The firm's short-run total cost function is C = 200 + 0.2q2, so its marginal cost is MC = 0.4q.
A. What is the firm's profit-maximizing quantity? (6 mks)
B. Is the firm earning a profit? (7 mks)
C. What is the short-run shutdown price? (7 mks)
1
Expert's answer
2016-12-05T06:28:14-0500
A. P=MC
0.4q=10
Q=25
B. Profit=TR-TC=10*25-200+0.2*10^2=250-220=30
C. 220=25*X
X=220/25=8.8